Reporting proposals for primary areas activity and associated projects - proposal for report on MEDICIS operation / Build. 179 coordination
b.179
New gas storage arriving week 6
RAMSES-CROME transition
PMIIS801 and PMIIS802 detectors removed yesterday
Preventative maintenance on UPS tomorrow (transparent to b.179)
Wed-Fri installation of access control for IBDRS
access to R-001
Safety valve for MEDICIS gas point and R-021 GB removed for maintenance
should be returned either today or tomorrow
vac team inspection of safety valves this afternoon
Decontamination shower not useable for next 3 days due to IBDRS work
MEDICIS
Monorail tests happened this morning especially link to the rescue shuttle. The rescue shuttle is not operational (and has never been) and has been removed.
Trajectories test will be done this afternoon.
Tomorrow Christophe will integrate the new interlock system into the rack
Two visit from CERN Society and fundation next week (21 and 22Jan), and the visit of the new collaborator from OSFP on 20 Jan organized by Fabio.
3
Known activities in the hall (EP side) - Proposal for coordination / reporting
Speaker:
Hanne Heylen(CERN)
Upcoming activities
Wisard offline runs (26/1-8/2)
Flammable gas lines - mostly done (no major works left). Installation of cage outside in week 5 (26/1).
4
Known activities in the hall (infrastructure / machine side) / Proposal for coordination and reporting
Speaker:
Erwin Siesling(CERN)
5
Discussion on organization of technical feedbacks session for ISOLDE systems
6
Review of invitation list / minutes distribution list